OpenStudy (anonymous):

square root question i am at the end of a problem and i have n1^2 = n2^2 so taking the square root shouldn't i get +or- n1 = = or - n2

OpenStudy (anonymous):

\[n _{1}^2 = n _{2}^2\] better formatting

OpenStudy (anonymous):

so taking the square root i should get \[\pm n _{1} = \pm n _{2}\]

OpenStudy (anonymous):

but the correct answer is \[n _{1} = \pm n _{2}\]

OpenStudy (anonymous):

what was the original problem

OpenStudy (anonymous):

proving a function is one to one. its a discrete math problem but for some reason i seem to have forgot how to do algebra

OpenStudy (anonymous):

prove n^2 - 1 is one to one, onto, or both is the original problem

OpenStudy (anonymous):

is n positive integers?

OpenStudy (anonymous):

the problem says the set of all integers for both domain and codomain so positive and negative

OpenStudy (anonymous):

f(n) = n^2 -1 , is not one to one f(3) = f(-3) but 3 \( \neq \) -3

OpenStudy (anonymous):

thats what i thought but was trying to prove it abstractly. but i guess i could just show one counter example and that would be proof enough it is not one to one

OpenStudy (anonymous):

so the problem to prove it directly \( n_1^2 -1 = n_2^2 -1 \) \( n_1^2 = n_2^2 \) you can't get n_1 = n_2 from this

OpenStudy (anonymous):

the our teacher did it was to assume that it is one to one and try and prove it or prove otherwise. We basically take the equation and set it equal to itself with the variable term being n1, n2 etc. which was weird to me

OpenStudy (anonymous):

\( n_1^2 -1 = n_2^2 -1\) \( n_1^2 = n_2^2 \) you have to stop here. you cannot deduce that \( n_1 = n_2 \) In general any counterexample works as follows: \(n_1 = -n_2 \) and \( n_1 \neq 0 \) then \( n_1^2 = (-n_2)^2 = n_2^2 \) but \( n_1 \neq n_2 \)

OpenStudy (anonymous):

oh, i see. i was not getting the algebra step you did. n1 = -n2 because its square they are equal and have multiple "x" values mapping to y values

OpenStudy (anonymous):

right, squaring undoes the negative (3)^2 = (-3)^2 but 3 \( \neq \) -3

OpenStudy (anonymous):

you can do a 'direct' one to one proof with a cubic function

OpenStudy (anonymous):

show that f(n) = n^3 -1 is one to one

OpenStudy (fibonaccichick666):

I may be able to help here, use a and b to make your life easier though

OpenStudy (anonymous):

also show that your function is not onto either hint use counterexample

OpenStudy (anonymous):

ok so \[n _{1}^3 = n _{2}^3\] then cube root them n1 = -n2 so for example f(3) and f(-3) 26 does not = -26

OpenStudy (fibonaccichick666):

If 1-1 \(f(a)\not=f(b)\) so for the squared one, let a be an integer \[f(a)=a^2+1\] \[f(-a)=(-a)^2+1=a^2+1\] Therefore f(a)=f(-a) and f is not 1-1

OpenStudy (anonymous):

so it would be one to one for the n^3-1

OpenStudy (fibonaccichick666):

yep

OpenStudy (fibonaccichick666):

thinking of a graph helps there. Does it pass the horizontal line test?

OpenStudy (fibonaccichick666):

for onto, you show every element has an inverse essentially.

OpenStudy (anonymous):

a graph does help i knw in my head that x^3 would be one to one so i just had to to a small proof

OpenStudy (anonymous):

yes it does pass

OpenStudy (fibonaccichick666):

so you use f(a) and f(b) for the n^3 one, just like I did for n^2 but in this instance you will come to the conclusion that they only equal if a=b which you stipulate is not possible at the beginning

OpenStudy (fibonaccichick666):

so, let \(a\not=b\) f(a)=f(b)...a=b therefore contradiction

OpenStudy (fibonaccichick666):

try filling in the rest

OpenStudy (fibonaccichick666):

oh and make sure you say a,b elem Z to start

OpenStudy (fibonaccichick666):

forgot that bit...always had a point taken off for it

OpenStudy (anonymous):

ok so y = n^3 -1 then y = \[\sqrt[3]{n^3 -1}\]

OpenStudy (fibonaccichick666):

wait,

OpenStudy (fibonaccichick666):

are you doing 1-1 or onto?

OpenStudy (anonymous):

i was going to do onto

OpenStudy (anonymous):

if you cube root a cube, you get \(n _{1}^3 = n _{2}^3 \\ \implies \\ \sqrt[3] {n _{1}^3} = \sqrt[3]{n _{2}^3} \)

OpenStudy (anonymous):

\( n_1 = n_2\)

OpenStudy (fibonaccichick666):

ok, can we just use a,b,c... much less confusing

OpenStudy (fibonaccichick666):

so for onto, I can do the square again

OpenStudy (fibonaccichick666):

so let a be real(in order to be onto, we need this to be a function onto the reals) then |a-1| is real, and sqrt|a-1| is real and\( f(\sqrt{|a-1|})=(\sqrt{|a-1|})^2+1=a-1=+/-a\) and +/-a is real, therefore there exists an y in R st f(x)=y for every x.

OpenStudy (fibonaccichick666):

onto the positive reals* abs. a-1 is +R

OpenStudy (fibonaccichick666):

well, to that effect. I messed something up in there

OpenStudy (anonymous):

ok

OpenStudy (fibonaccichick666):

X^3+1 would be easier. since no potential complex numbers x^3+1 (x-1)^1/3

OpenStudy (anonymous):

oh for the n^2 -1 the problem was restricted to integers. So for onto could i just do y=x^2 - 1 y=sqrt(x^2-1) y(1) = sqrt(2) therefore since i get a number not in the original set of integers i could say its not onto?

OpenStudy (anonymous):

but i didnt realize that until i saw what you did

OpenStudy (anonymous):

if n = -4 which is an integer there is no value such that f(n) = -4

OpenStudy (anonymous):

i knew it was restricted to integers but i didnt realize i could use that as the reason its not onto

OpenStudy (fibonaccichick666):

yea, pick a number and make it a counter ex

OpenStudy (anonymous):

ok so for onto there are many integers that f(n) is not in the domain

OpenStudy (fibonaccichick666):

yep, all you need is 1 value that is in the range that has no x to get you there

OpenStudy (anonymous):

ok. I think i get it now. its easy to show its not one to one and not onto. just find a counter example. proving it is one to one or onto is going to be a little harder.

OpenStudy (fibonaccichick666):

... but the x^2 is onto. The range of the function x^2+1 is [1,infty)

OpenStudy (fibonaccichick666):

sqrt 4 is 2,+-2

OpenStudy (fibonaccichick666):

because it isn't one to one, it can be onto.

OpenStudy (fibonaccichick666):

you just run into defining the range being necessary

OpenStudy (fibonaccichick666):

we are defining a positive domain so a sqrt has a real result

OpenStudy (anonymous):

wait. maybe i misunderstand onto. but it the problem says the domain and codomain are all integers then y=x^2 - 1 y=sqrt(x^2-1) y(3) = sqrt(3^2 -1) y=(sqrt(8)) therefore since i get a number not in the integers wouldnt it be not onto?

OpenStudy (anonymous):

also it might help remember for \( \large x \in \mathbb R \) , \( \large \sqrt{x^2}= |x| \) but \( \large \sqrt[3]{x^3}= x \) Using this if \(\large n_2^2 = n_1^2 \) if we square root both sides we get \(\large \sqrt{n_2^2 } = \sqrt{n_1^2} \iff |n_2| = |n_1| \) but you cannot remove the absolute value bars to get \( \large n_2 = n_1 \)

OpenStudy (anonymous):

similarly with even roots same deal. and odd roots

OpenStudy (fibonaccichick666):

you don't have an accurate inverse it's \[y=\sqrt{x-1}\]

OpenStudy (anonymous):

oh ok. but if x is 8 then i get sqrt(7) which is still not an integer. so its not it the set of integers specified by the problem?

OpenStudy (fibonaccichick666):

since the squared cancels out it doesn't matter if sqrt(x-1) is negative(which it cannot be given the definition of its domain(our prior range)

OpenStudy (fibonaccichick666):

you have to do f(y)

OpenStudy (fibonaccichick666):

not just compute y

OpenStudy (anonymous):

wait isnt the inverse y = \[\sqrt{x+1} \] and not y= \[\sqrt{x-1}\]

OpenStudy (fibonaccichick666):

f(x)=x^2+1 y=x^2+1 switch x=y^2+1 x-1=y^2 sqrt(x-1)=y

OpenStudy (anonymous):

im still not seeing how it is onto. i say its not because there are integers the domain that make the codomain not in the set of integers

OpenStudy (fibonaccichick666):

to be onto, you must be able to hit every element of the range

OpenStudy (fibonaccichick666):

here our function is Z->[1,infty)

OpenStudy (anonymous):

oh, ok. so even though there are numbers not in the range. there are numbers that can be found for all integers which is our range

OpenStudy (fibonaccichick666):

yea. That function is not onto the integers. It is onto a subset of the integers. So we are not going backwards from the whole integers, that wasn't our range. We are just taking our range and going back.

OpenStudy (fibonaccichick666):

This slideshow may help http://www.csee.umbc.edu/~stephens/203/PDF/7-3.pdf

OpenStudy (anonymous):

every function is onto its range, trivially :)

OpenStudy (anonymous):

not every function is onto its 'codomain'

OpenStudy (fibonaccichick666):

function implies surjective

OpenStudy (fibonaccichick666):

but you still have to learn to prove it

OpenStudy (fibonaccichick666):

it's part of proving something is a function.

OpenStudy (fibonaccichick666):

here is the thing though, if a problem were to say f(x):R-->R then it would have to span the whole reals. we weren't given something like that here, so we assume the range makes sense

OpenStudy (anonymous):

ok. slide 7.3.6 is what i need.

OpenStudy (anonymous):

im looking at the slides so i might be gone a few minutes

OpenStudy (fibonaccichick666):

is cool

OpenStudy (anonymous):

but i really appreciate both of you guys helping. the one to one isnt that bad. the onto i think is where i need practice.

OpenStudy (fibonaccichick666):

it's not as terrible as it feels. Do a checklist: what is my range? Do I hit my whole range? yes?-->onto no?--->not onto

OpenStudy (fibonaccichick666):

eg. f(x):Z-->Z f(x)=x^2+4 What's my range? Z Do I hit every point? No Not onto. e.g. Let y=0 x^2=-4-->x=2i which is not in Z.

OpenStudy (anonymous):

you call the 'range' what i call the 'codomain'

OpenStudy (fibonaccichick666):

they are the same thing, but we didn't have one defined before

OpenStudy (fibonaccichick666):

ok, maybe not logically the exact same thing :P

OpenStudy (anonymous):

a function f: N -> R , N is domain and R is codomain and f(N) i call range. some people call it the image .

OpenStudy (fibonaccichick666):

I call it image. But to be honest, it's semantics to me. I learned range first it stuck.

OpenStudy (anonymous):

got it. Im going to use the letters in my problem f(n) = n^2 -1 Let y be in Y n^2 - 1 = y n^2 = y+1 n= sqrt(y +1) f(n) = f(sqrt(y+1)) = (sqrt (y+1))^2 -1 = y +1 -1 = y therefore it exists

OpenStudy (anonymous):

sorry for the bad formatting but im slow using the equation and inserting them

OpenStudy (fibonaccichick666):

ok, now, leave out the middle bit. Just state a y value

OpenStudy (fibonaccichick666):

the actually finding the inverse is scratch work

OpenStudy (anonymous):

i guess 2, 3 or 4 any really

OpenStudy (anonymous):

oh well dang.

OpenStudy (fibonaccichick666):

oh, no I mean just claim y=inverse essentially